# Solvable families of random block tridiagonal matrices

### v1

@@ 3.4 A reduction and a representation
The point is that by Lemma 32 below we have the following. Suppose that $b_{i},1\leq i\leq 8$ are fixed nonnegative integers with $b_{i}=b_{9-i}$ and $\sum_{i=1}^{4}b_{i}=n$. Let ${\mathcal{P}}_{\underline{b}}$ be the collection of partitions ¯$\underline{\mathcal{B}}=(\mathcal{B}_{1},\dots,\mathcal{B}_{8})$ of $[\![2n]\!]$ with $|\mathcal{B}_{i}|=b_{i},1\leq i\leq 8$. Then
¯$\sum_{\underline{\mathcal{B}}\in{\mathcal{P}}_{\underline{b}}}\prod_{i=1}^{4}C_{\mathcal{B}_{i},\mathcal{B}_{9-i}}^{2}=\binom{n}{b_{1},b_{2},b_{3},b_{4}}\sum_{\mathcal{A}\subset[\![2n]\!],|\mathcal{A}|=n}C_{\mathcal{A},\mathcal{A}^{\prime}}^{2}.$ (91)
@@ 6 Determinantal identities
Let $n\geq 2$, and suppose that $b_{i},1\leq i\leq 8$ are fixed nonnegative integers with $b_{i}=b_{9-i}$ and $\sum_{i=1}^{4}b_{i}=n$. Let ${\mathcal{P}}_{\underline{b}}$ be the set of partitions ¯$\underline{\mathcal{B}}=(\mathcal{B}_{1},\dots,\mathcal{B}_{8})$ of $[\![2n]\!]$ with $|\mathcal{B}_{i}|=b_{i},1\leq i\leq 8$. Then
¯$\sum_{\underline{\mathcal{B}}\in{\mathcal{P}}_{\underline{b}}}\prod_{j=1}^{4}C_{\mathcal{B}_{j},\mathcal{B}_{9-j}}^{2}=\binom{n}{b_{1},b_{2},b_{3},b_{4}}\sum_{\mathcal{A}\subset[\![2n]\!],|\mathcal{A}|=n}C_{\mathcal{A},\mathcal{A}^{\prime}}^{2}.$ (95)
@@ 6.2 Proofs of Lemmas 32 and 33
Here $\Pi_{2n}$ refers to the set of matchings on $[\![2n]\!]$, i.e, partitions $(\alpha_{1,1},\alpha_{1,2}),\dots(\alpha_{n,1},\alpha_{n,2})$ of $[\![2n]\!]$ with $\alpha_{i,1}<\alpha_{i,2}$ and $\alpha_{1,1}<\cdots<\alpha_{n,1}$. In the Pfaffian formula, $\pi_{\alpha}$ denotes the corresponding permutation $(\alpha_{1,1},\alpha_{1,2},\alpha_{2,1},\alpha_{2,2}\dots\alpha_{n,1},\alpha_{n,2})$.
@@
Fix a length $k-1$ cycle $(b_{1},\dots,b_{k-1})$ on $\{1,\dots,k-1\}$, and consider all length $k$ cycles that we get by inserting $k$ at some point. Then by (109) the contribution of all of these cycles in the sum in (107) (using $b_{k}=b_{1}$) is
@@
Consider the mapping that takes a particular choice of $a_{1},\dots,a_{n},a_{1}^{\prime},\dots,a_{n}^{\prime}$ and $\sigma\in S_{n}$ into the matching of $[\![2n]\!]$ that matches $a_{i}$ with $a_{\sigma(i)}^{\prime}$ for $1\leq i\leq n$. Each particular matching $\alpha\in\Pi_{2n}$ shows up exactly $2^{n}(n!)^{2}$ times as the result of this mapping, which proves (113).
@@
For a particular pair $\sigma,\tau$ consider the permutation on $[\![2n]\!]=\{a_{1},\dots,a_{n},a_{1}^{\prime},\dots,a_{n}^{\prime}\}$ that takes $a_{i}$ to $a_{\sigma_{i}}^{\prime}$ and $a_{i}^{\prime}$ to $a_{\tau^{-1}(i)}$. This permutation has even cycles, and because $\sigma\neq\tau$ it must have at least one cycle of length at least 4. Let one of these cycles be
@@
The sum on the left can be rewritten as a double sum where we first sum over all possible assignments of the values of $a_{i_{1}},a^{\prime}_{j_{1}},a_{i_{2}},a^{\prime}_{j_{2}}\dots,a_{i_{k}},a^{\prime}_{j_{k}}$, and then in the second sum we sum over the remaining $2n-2k$ variables. Factoring out the terms corresponding to the indices not in $\mathcal{B}$, the inner sum becomes
@@
where the sum is over all the permutations of $[\![2n]\!]$ that have no fixed elements. Because of Lemma 38, the contribution of the permutations that have a cycle of length at least 3 cancels out. The remaining terms correspond to the permutations that only have 2-cycles, that is, the permutations that corresponding to perfect matchings of $[\![2n]\!]$. For such a permutation $\sigma$ we have
@@
¯$\sum_{\underline{\mathcal{B}}\in{\mathcal{P}}_{\underline{b}}}\prod_{i=1}^{4}C_{\mathcal{B}_{i},\mathcal{B}_{9-i}^{2}}=\frac{1}{(b_{1}!b_{2}!b_{3}!b_{4}!)^{2}}\sum_{\sigma\in S_{2n}}\prod_{j=1}^{4}C^{2}_{\mathcal{B}_{\sigma,j},\mathcal{B}^{\prime}_{\sigma,j}}.$ (118)
@@
where $S(\mathcal{D}_{j})$ is the set of permutations of $\mathcal{D}_{j}$. We can represent $\eta_{1},\dots,\eta_{4}$ as a single permutation of $[\![n]\!]$ that preserves $\mathcal{D}_{1},\mathcal{D}_{2},\mathcal{D}_{3},\mathcal{D}_{4}$. Denoting the set of these permutations $S_{b_{1},b_{2},b_{3},b_{4}}$ we get
¯$\sum_{\underline{\mathcal{B}}\in{\mathcal{P}}_{\underline{b}}}\prod_{j=1}^{4}C_{\mathcal{B}_{j},\mathcal{B}_{9-j}}^{2}=\frac{1}{(b_{1}!b_{2}!b_{3}!b_{4}!)^{2}}\sum_{\sigma\in S_{2n}}\sum_{\eta,\tilde{\eta}\in S_{b_{1},b_{2},b_{3},b_{4}}}\operatorname{sgn}(\eta)\operatorname{sgn}(\tilde{\eta})\prod_{i=1}^{n}\frac{1}{z_{i}-z^{\prime}_{\eta(i)}}\cdot\frac{1}{z_{i}-z^{\prime}_{\tilde{\eta}(i)}}.$ (120)
@@
where $(\alpha_{1,1},\alpha_{1,2}),\dots(\alpha_{n,1},\alpha_{n,2})$ is a matching of $[\![2n]\!]$, i.e. an element of $\Pi_{2n}$. The term corresponding to a particular matching $\alpha\in\Pi_{2n}$ shows up in the sum in (121) exactly $2^{n}n!b_{1}!b_{2}!b_{3}!b_{4}!$ times. Indeed:
@@
To prove this statement, we first note that the pair $\eta,\tilde{\eta}$ generates a permutation of $[\![2n]\!]$ with $i\to n+\eta(i)$, $n+i\to\tilde{\eta}^{-1}(i)$, $1\leq i\leq n$. This permutation has only even length cycles, and since $\eta\neq\tilde{\eta}$, it has at least one cycle with length $2\ell\geq 4$. Let one of these cycles be
@@
=$\mathcal{F}=\{i_{1},\dots,i_{\ell},n+j_{1},\dots,n+j_{\ell}\}.$
@@ 7.1 The ${\tt{H}}_{\beta,n}(r,s)$ soft edge
Denote by $\mu_{\gamma}$ the semi-circle distribution with density $\frac{1}{\pi\gamma}\sqrt{(4\gamma-\lambda^{2})_{+}}$. For any integer $r$ any $s>1-r$, the empirical spectral measure $\frac{1}{rn}\sum_{i=1}^{rn}\delta_{\lambda_{i}}$ of eigenvalues of $\mathtt{H}_{\beta,n}(r,s)$ converges weakly in law to $\mu_{\gamma}$ with $\gamma=\frac{r+s}{r}$.
@@
$\mathbf{H}_{n}$ $=\gamma^{-1/2}(rn)^{1/6}(2\sqrt{(r+s)n}{I}_{rn}-\mathbf{T}_{n})$ (124)
@@
Said another way, for integers $i\in[0,r-1]$ and $k\in[0,n]$, $v_{i+kr}$ is identified with $v_{i}(k/m_{n})$ With this embedding in mind, write $\mathbf{H}_{n}=m_{n}^{2}\Delta_{n}+m_{n}{\mathbf{V}}_{n}$. Here $\Delta_{n}$ denotes discrete Laplacian in the underlying basis: it has diagonal blocks $2{I}_{r}$ and off-diagonal blocks $-{I}_{r}$. The $\mathbf{V}_{n}$ is then considered to define a matrix-valued potential. The main technical result of [2] provides criteria for the convergence of the spectrum of $\mathbf{H}_{n}$ (in law) in terms of the convergence of the integrated components of $\mathbf{V}_{n}$. A version of this, tailored to the particular setting considered here, is the following.
@@
${\bf{Y}}_{n,1}(x)=\sum_{i=1}^{[m_{n}x]}\left(-\frac{1}{\sqrt{\gamma m_{n}}}\mathbf{A}_{n,i}\right)\quad{\bf{Y}}_{n,2}(x)=\sum_{i=1}^{[m_{n}x]}\left(m_{n}^{2}\mathbf{I}_{r}-\frac{1}{\sqrt{\gamma m_{n}}}\mathbf{B}_{n,i}\right).$ (126)
@@
(ii) For $j=1,2$ there is a decomposition of the increments,
$(\Delta{\bf{Y}})_{n,j}=\boldsymbol{\eta}_{n,j}+(\Delta{\boldsymbol{\omega}})_{n,j},$ (128)
where: (1) the $\boldsymbol{\eta}_{n,j,i}$ are diagonal and are bounded entry-wise as in
@@
[…]
@@
for all $|x-y|\leq 1$ and $\kappa_{n}\geq 1$ a sequence of tight random constants.
@@
One may then verify that points $x^{\prime}$ where $\det F$ vanishes correspond to $P(x)$ possessing an eigenvalue $p(x)$ that explodes to $\infty$- ∞ as $x\rightarrow x^{\prime}$. This is exactly the content of Theorem 3: the stochastic differential equation (5), as can be derived from (132) by an application of Itô’s Lemma, describes the evolution of the eigenvalues $(p_{1}(x),\dots,p_{r}(x))$ of $P(x)$, continued through explosion times to $\infty$- ∞. (Again, see [2] for additional details.)
@@ 7.2 Hard edge for ${\tt{W}}_{\beta,n,n+a}(r,s)$
[…]
@@
The results of [21] imply that $\mathcal{L}_{\beta,\gamma}^{-1}$ is Hilbert-Schmidt for any $\gamma>0$ and $a>-1$, and the main convergence result there can be summarized, in the spirit of Proposition 40, as follows.
@@
Note that while the embedding of the matrix $\mathbf{L}_{n}^{-1}$ naturally takes place as an operator on $[0,1]$, the advertised limit $\mathcal{L}^{-1}_{\beta,\gamma}$ lives on $[0,\infty)$. This is just more convenient for eventual comparison to the soft edge operator $\mathcal{H}_{\beta,\gamma}$, and the kernel $\ell(x,y)$ is related to the limit of $\ell_{n}$, which is constructed from the limit objects defined in (i) above, by the simple change of variables $x\mapsto 1-e^{-rx}=\varphi(x)$. As one can readily check, $\mathbf{X}_{\varphi(x)}=\mathbf{Y}_{x}$ and $\frac{1}{r\sqrt{1-\varphi(x)}}d\varphi(x)=e^{-rx/2}dx$. This establishes Theorem 7
@@
acting on the positive half-line. This is well defined for all $\beta>0$ and $a>-1$.

### v2

@@ 3.4 A reduction and a representation
The point is that by Lemma 32 below we have the following. Suppose that $b_{i},1\leq i\leq 8$ are fixed nonnegative integers with $b_{i}=b_{9-i}$ and $\sum_{i=1}^{4}b_{i}=n$. Let ${\mathcal{P}}_{\underline{b}}$ be the collection of partitions $\underline{\mathcal{B}}=(\mathcal{B}_{1},\dots,\mathcal{B}_{8})$ of [[2n]][\![2n]\!] with $|\mathcal{B}_{i}|=b_{i},1\leq i\leq 8$. Then
$\sum_{\underline{\mathcal{B}}\in{\mathcal{P}}_{\underline{b}}}\prod_{i=1}^{4}C_{\mathcal{B}_{i},\mathcal{B}_{9-i}}^{2}=\binom{n}{b_{1},b_{2},b_{3},b_{4}}\sum_{\mathcal{A}\subset[\![2n]\!],|\mathcal{A}|=n}C_{\mathcal{A},\mathcal{A}^{\prime}}^{2}.$ (91)
@@ 6 Determinantal identities
Let $n\geq 2$, and suppose that $b_{i},1\leq i\leq 8$ are fixed nonnegative integers with $b_{i}=b_{9-i}$ and $\sum_{i=1}^{4}b_{i}=n$. Let ${\mathcal{P}}_{\underline{b}}$ be the set of partitions $\underline{\mathcal{B}}=(\mathcal{B}_{1},\dots,\mathcal{B}_{8})$ of [[2n]][\![2n]\!] with $|\mathcal{B}_{i}|=b_{i},1\leq i\leq 8$. Then
$\sum_{\underline{\mathcal{B}}\in{\mathcal{P}}_{\underline{b}}}\prod_{j=1}^{4}C_{\mathcal{B}_{j},\mathcal{B}_{9-j}}^{2}=\binom{n}{b_{1},b_{2},b_{3},b_{4}}\sum_{\mathcal{A}\subset[\![2n]\!],|\mathcal{A}|=n}C_{\mathcal{A},\mathcal{A}^{\prime}}^{2}.$ (95)
@@ 6.2 Proofs of Lemmas 32 and 33
Here $\Pi_{2n}$ refers to the set of matchings on [[2n]][\![2n]\!], i.e, partitions $(\alpha_{1,1},\alpha_{1,2}),\dots(\alpha_{n,1},\alpha_{n,2})$ of [[2n]][\![2n]\!] with $\alpha_{i,1}<\alpha_{i,2}$ and $\alpha_{1,1}<\cdots<\alpha_{n,1}$. In the Pfaffian formula, $\pi_{\alpha}$ denotes the corresponding permutation $(\alpha_{1,1},\alpha_{1,2},\alpha_{2,1},\alpha_{2,2}\dots\alpha_{n,1},\alpha_{n,2})$.
@@
Fix a length k1k-1 cycle $(b_{1},\dots,b_{k-1})$ on $\{1,\dots,k-1\}$, and consider all length $k$ cycles that we get by inserting $k$ at some point. Then by (109) the contribution of all of these cycles in the sum in (107) (using $b_{k}=b_{1}$) is
@@
Consider the mapping that takes a particular choice of $a_{1},\dots,a_{n},a_{1}^{\prime},\dots,a_{n}^{\prime}$ and $\sigma\in S_{n}$ into the matching of [[2n]][\![2n]\!] that matches $a_{i}$ with $a_{\sigma(i)}^{\prime}$ for $1\leq i\leq n$. Each particular matching $\alpha\in\Pi_{2n}$ shows up exactly $2^{n}(n!)^{2}$ times as the result of this mapping, which proves (113).
@@
For a particular pair $\sigma,\tau$ consider the permutation on $2n]\!]=\{a_{1},\dots,a_{n},a_{1}^{\prime},\dots,a_{n}^{\prime}\}$ that takes $a_{i}$ to $a_{\sigma_{i}}^{\prime}$ and $a_{i}^{\prime}$ to $a_{\tau^{-1}(i)}$. This permutation has even cycles, and because $\sigma\neq\tau$ it must have at least one cycle of length at least 4. Let one of these cycles be
@@
The sum on the left can be rewritten as a double sum where we first sum over all possible assignments of the values of $a_{i_{1}},a^{\prime}_{j_{1}},a_{i_{2}},a^{\prime}_{j_{2}}\dots,a_{i_{k}},a^{\prime}_{j_{k}}$, and then in the second sum we sum over the remaining 2n2k2n-2k variables. Factoring out the terms corresponding to the indices not in $\mathcal{B}$, the inner sum becomes
@@
where the sum is over all the permutations of [[2n]][\![2n]\!] that have no fixed elements. Because of Lemma 38, the contribution of the permutations that have a cycle of length at least 3 cancels out. The remaining terms correspond to the permutations that only have 2-cycles, that is, the permutations that corresponding to perfect matchings of [[2n]][\![2n]\!]. For such a permutation $\sigma$ we have
@@
$\sum_{\underline{\mathcal{B}}\in{\mathcal{P}}_{\underline{b}}}\prod_{i=1}^{4}C_{\mathcal{B}_{i},\mathcal{B}_{9-i}^{2}}=\frac{1}{(b_{1}!b_{2}!b_{3}!b_{4}!)^{2}}\sum_{\sigma\in S_{2n}}\prod_{j=1}^{4}C^{2}_{\mathcal{B}_{\sigma,j},\mathcal{B}^{\prime}_{\sigma,j}}.$ (118)
@@
where $S(\mathcal{D}_{j})$ is the set of permutations of $\mathcal{D}_{j}$. We can represent $\eta_{1},\dots,\eta_{4}$ as a single permutation of [[n]][\![n]\!] that preserves $\mathcal{D}_{1},\mathcal{D}_{2},\mathcal{D}_{3},\mathcal{D}_{4}$. Denoting the set of these permutations $S_{b_{1},b_{2},b_{3},b_{4}}$ we get
$\sum_{\underline{\mathcal{B}}\in{\mathcal{P}}_{\underline{b}}}\prod_{j=1}^{4}C_{\mathcal{B}_{j},\mathcal{B}_{9-j}}^{2}=\frac{1}{(b_{1}!b_{2}!b_{3}!b_{4}!)^{2}}\sum_{\sigma\in S_{2n}}\sum_{\eta,\tilde{\eta}\in S_{b_{1},b_{2},b_{3},b_{4}}}\operatorname{sgn}(\eta)\operatorname{sgn}(\tilde{\eta})\prod_{i=1}^{n}\frac{1}{z_{i}-z^{\prime}_{\eta(i)}}\cdot\frac{1}{z_{i}-z^{\prime}_{\tilde{\eta}(i)}}.$ (120)
@@
where $(\alpha_{1,1},\alpha_{1,2}),\dots(\alpha_{n,1},\alpha_{n,2})$ is a matching of [[2n]][\![2n]\!], i.e. an element of $\Pi_{2n}$. The term corresponding to a particular matching $\alpha\in\Pi_{2n}$ shows up in the sum in (121) exactly $2^{n}n!b_{1}!b_{2}!b_{3}!b_{4}!$ times. Indeed:
@@
To prove this statement, we first note that the pair $\eta,\tilde{\eta}$ generates a permutation of [[2n]][\![2n]\!] with $i\to n+\eta(i)$, $n+i\to\tilde{\eta}^{-1}(i)$, $1\leq i\leq n$. This permutation has only even length cycles, and since $\eta\neq\tilde{\eta}$, it has at least one cycle with length $2\ell\geq 4$. Let one of these cycles be
@@
$\mathcal{F}=\{i_{1},\dots,i_{\ell},n+j_{1},\dots,n+j_{\ell}\}.$
@@ 7.1 The ${\tt{H}}_{\beta,n}(r,s)$ soft edge
Denote by $\mu_{\gamma}$ the semi-circle distribution with density $\frac{1}{\pi\gamma}\sqrt{(4\gamma-\lambda^{2})_{+}}$. For any integer $r$ any s>1rs>1-r, the empirical spectral measure $\frac{1}{rn}\sum_{i=1}^{rn}\delta_{\lambda_{i}}$ of eigenvalues of $\mathtt{H}_{\beta,n}(r,s)$ converges weakly in law to $\mu_{\gamma}$ with $\gamma=\frac{r+s}{r}$.
@@
$\mathbf{H}_{n}$ $=\gamma^{-1/2}(rn)^{1/6}(2\sqrt{(r+s)n}\mathbf{I}_{rn}-\mathbf{T}_{n})$ (124)
@@
Said another way, for integers $i\in[0,r-1]$ and $k\in[0,n]$, $v_{i+kr}$ is identified with $v_{i}(k/m_{n})$ With this embedding in mind, write $\mathbf{H}_{n}=m_{n}^{2}\Delta_{n}+m_{n}{\mathbf{V}}_{n}$. Here $\Delta_{n}$ denotes discrete Laplacian in the underlying basis: it has diagonal blocks $2{I}_{r}$ and off-diagonal blocks ${I}_{r}$. The $\mathbf{V}_{n}$ is then considered to define a matrix-valued potential. The main technical result of [2] provides criteria for the convergence of the spectrum of $\mathbf{H}_{n}$ (in law) in terms of the convergence of the integrated components of $\mathbf{V}_{n}$. A version of this, tailored to the particular setting considered here, is the following.
@@
${\bf{Y}}_{n,1}(x)=\sum_{i=1}^{[m_{n}x]}\left(-\frac{1}{\sqrt{\gamma m_{n}}}\mathbf{A}_{n,i}\right),\quad{\bf{Y}}_{n,2}(x)=\sum_{i=1}^{[m_{n}x]}\left(m_{n}^{2}\mathbf{I}_{r}-\frac{1}{\sqrt{\gamma m_{n}}}\mathbf{B}_{n,i}\right).$ (126)
@@
(ii) For $j=1,2$, defining the increments via ${\bf{Y}}_{n,j}(x):=\frac{1}{\sqrt{\gamma m_{n}}}\sum_{i=1}^{[m_{n}x]}(\Delta{\bf{Y}})_{n,j,i}$ there is a decomposition,
$(\Delta{\bf{Y}})_{n,j,i}=\boldsymbol{\eta}_{n,j,i}+(\Delta{\boldsymbol{\omega}})_{n,j,i}$ (128)
such that the following holds. The $\boldsymbol{\eta}_{n,j,i}$ are diagonal and $\boldsymbol{\eta}_{n,j}(x)=\boldsymbol{\eta}_{n,j,[m_{n}x]}$ satisfy
@@
for some $\kappa\geq 1$. Also, with $\boldsymbol{\omega}_{n,j}(x)=\sum_{i=1}^{[m_{n}x]}(\Delta{\boldsymbol{\omega}})_{n,j,i}$, there exists an $\epsilon>0$ so that
@@
for all $|x-y|\leq 1$. Here $\|\cdot\|$ denotes the spectral norm, and $\kappa_{n}\geq 1$ a sequence of tight random constants.
@@
One may then verify that points $x^{\prime}$ where $\det F$ vanishes correspond to $P(x)$ possessing an eigenvalue $p(x)$ that explodes to -$\infty$ as $x\rightarrow x^{\prime}$. This is exactly the content of Theorem 3: the stochastic differential equation (5), as can be derived from (132) by an application of Itô’s Lemma, describes the evolution of the eigenvalues $(p_{1}(x),\dots,p_{r}(x))$ of $P(x)$, continued through explosion times to -$\infty$. (Again, see [2] for additional details.)
@@ 7.2 Hard edge for ${\tt{W}}_{\beta,n,n+a}(r,s)$
$\mathcal{L}_{\beta,\gamma}^{-1}f(x)=\int_{x}^{\infty}{\ell}(x,y)f(y)dy,\quad{\ell}(x,y)=e^{-rx/2}\mathbf{Y}_{x}^{-1}\mathbf{Y}_{y}.$ (134)
@@
The results of [21] imply that $\mathcal{L}_{\beta,\gamma}^{-1}$ is Hilbert-Schmidt for any $\gamma>0$ and a>1a>-1, and the main convergence result there can be summarized, in the spirit of Proposition 40, as follows.
@@
Note that while the embedding of the matrix $\mathbf{L}_{n}^{-1}$ naturally takes place as an operator on $[0,1]$, the advertised limit $\mathcal{L}^{-1}_{\beta,\gamma}$ lives on $0,\infty)$. This is just more convenient for eventual comparison to the soft edge operator $\mathcal{H}_{\beta,\gamma}$, and the kernel $\ell(x,y)$ is related to the limit of $\ell_{n}$, which is constructed from the limit objects defined in (i) above, by the simple change of variables $x\mapsto 1-e^{-rx}=\varphi(x)$. As one can readily check, $\mathbf{X}_{\varphi(x)}=\mathbf{Y}_{x}$ and $\frac{1}{r\sqrt{1-\varphi(x)}}d\varphi(x)=e^{-rx/2}dx$. This establishes Theorem 7
@@
acting on the positive half-line. This is well defined for all $\beta>0$ and a>1a>-1.